# On Supportedness in Multi-Objective Combinatorial Optimization

### v1

@@ 3 Supported and Weakly Supported Non-dominated Points in MOCO problems
Figure 4: Illustration of Example 3.11 highlighting the points $y^{1},y^{2}$ the points that contains the points that are both in $\mathcal{Y}_{SNF}$ and $\mathcal{Y}_{S\lambda}$ (bold curve), which together with $y^{1}$ and $y^{2}$ gives the non-dominated frontier.
@@
$\mathcal{Y}_{S\lambda}\subseteq\mathcal{Y}_{SNF}$ follows directly from the proof of 3.6. Furthermore, as demonstrated in Example 3.11, there exist instances where $\mathcal{Y}_{S\lambda}\subset\mathcal{Y}_{SNF}$. ∎
The distinction between supported and weakly supported is sufficient for MOCO problems. However, in the context of general MOO, it may be worthwhile to introduce a finer classification, distinguishing between properly supported, supported, and weakly supported non-dominated points. While this work primarily focuses on the supportedness in MOCO problems, a comprehensive analysis of supportedness in general MOO problems remains an important area for future research. Notably, recent advancements in this direction have been discussed in Chlumsky-Harttmann and Schöbel, (2024).
@@ 4 Conclusion
Using these refined definitions, the following characterizations for MOCO problems hold: While the weakly supported non-dominated points lie on the boundary of the upper image, the supported non-dominated points lie only on the non-dominated frontier, i. e., only on maximally non-dominated faces, and the unsupported non-dominated points lie in the interior of the upper image. However, these characterizations do not extend directly to general MOO problems. Hence, future research should aim to extend the characterization of supportedness beyond MOCO problems to general MOO contexts, addressing the gaps highlighted by this study. An approach towards a categorization of supportedness in the context of general multi-objective optimization has been presented at the EURO conference 2024 (Chlumsky-Harttmann and Schöbel,, 2024).

### v2

@@ 2 Supported and Weakly Supported Non-dominated Points in MOCO problems
Figure 5: Illustration of Example˜2.10 highlighting the points $y^{1},y^{2}$ the points that contains the points that are both in $\mathcal{Y}_{SNF}$ and $\mathcal{Y}_{S\lambda}$ (bold curve), which together with $y^{1}$ and $y^{2}$ gives the non-dominated frontier.
@@
$\mathcal{Y}_{S\lambda}\subseteq\mathcal{Y}_{SNF}$ follows directly from the proof of ˜2.5. Furthermore, as demonstrated in Example˜2.10, there exist instances where $\mathcal{Y}_{S\lambda}\subset\mathcal{Y}_{SNF}$. ∎
The distinction between supported and weakly supported is sufficient for MOCO and MOILP problems. However, in the context of general MOO, it may be worthwhile to introduce a finer classification, distinguishing between properly supported, supported, and weakly supported non-dominated points. While this work primarily focuses on the supportedness in MOCO problems, a comprehensive analysis of supportedness in general MOO problems remains an important area for future research. Notably, recent advancements in this direction have been discussed in Chlumsky-Harttmann, (2025).
@@ 3 Conclusion
Using these refined definitions, the following characterizations for MOCO problems hold: While the weakly supported non-dominated points lie on the boundary of the upper image, the supported non-dominated points lie only on the non-dominated frontier, i. e., only on maximally non-dominated faces, and the unsupported non-dominated points lie in the interior of the upper image. However, these characterizations do not extend directly to general MOO problems. Hence, future research should aim to extend the characterization of supportedness beyond MOCO problems to general MOO contexts, addressing the gaps highlighted by this study. An approach towards a categorization of supportedness in the context of general multi-objective optimization has been discussed in Chlumsky-Harttmann, (2025).